A perfectly competitive firm facing the following TVC function: TVC= 150Q - 20Q² + Q³, where q= quantity.

Below what price should the firm shutdown it's operations


1
Expert's answer
2021-04-05T11:03:26-0400
"solution"

The firm should shut down operations if price falls below "\$50."

Explanation

The shut-down point of a company is a point whereby the marginal cost ("MC" ) of a firm is equal to the average variable cost ("AVC" ).

Marginal cost of a firm is a derivative of the total cost with respect to the quantity.

Therefore;


"MC=\frac{dTVC}{dQ}\\

 MC =\frac{150Q - 20Q^2 + Q^3}{Q}\\

 MC =150 -40Q +3Q^2"


According to profit maximization theory, "P=MC" , thus

"P= 150 -40Q +3Q^2"

Average variable equation is given by;

"AVC = \frac{TVC}{Q}\\

 AVC =\frac{150Q – 20Q2 +Q3)} { Q}\\

 AVC = 150 -20Q +Q2"

Equate the two equations, that is "MC" and "AVC"

"MC=AVC"

 "150 -40Q +3Q^2 =150 -20Q +Q^2\\

2Q^2-20Q \ or\\

2Q (Q-10) =0"

Thus, "Q = 0" and "Q = 10"

substitute "Q=10" in the marginal cost equation

"MC=150 -40Q +3Q^2\\

MC=150-40(10) +3(100)\\

MC=150-400+300\\

MC=50"

Therefore, "P=\$50."

Also substitute "Q=0" in the marginal equation

"MC= 150 -40Q +3Q^2\\

MC=150-40(0) +0\\

MC=150"

Therefore, "P=\$150"

A firm will shut down operations when "P < AVC"

Thus, in this case the firm will shut down when the price falls below "\$50."